RC Meeting and Run Plan, Monday, February 25

From HallCWiki
Jump to: navigation, search

Main | Online | RC | Run Plan | Official Website | Redmine | Index | Nav

RC Meeting Information

RC Daily Meetings are at 4:30 PM, 2nd floor Counting House Meeting Room

  • Bluejeans remote connection information:
 To join the Meeting:
 To join via Room System:
 Video Conferencing System: bjn.vc -or-
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920

Useful Links

Current Status

  • Continued running J/Psi on kinematic setting 3 high-t over the weekend.
  • Will switch to phase-1 kinematic at 1:00 pm on Monday. This requires turning on the rotation controllers.

Current Run Plan

Notices in Effect for Shift Crew

  • Shift Leader is responsible for 50k replays. Even if the TO is actually doing them.
  • Target Operators: When the alarm handler goes white, it is disconnected from the IOC. Log it and include carlin@jlab.org every time. Acknowledge the warning. If it is still white and not coming back, restart the alarm handler. Do not permanently silence the alarm handler.
  • Missing Reference Times during the replay: check in the 50k. Check in full replay, if able. If it happens, call RC. Guidance to solve it is here: https://logbooks.jlab.org/entry/3654584

Current Issues

  • SHMS Q1 Manually set (need to fix comm issues later). Read-back current masked in alarm handler.
  • Rotation Controllers turned back on this morning.
  • Concerning target alarms (loss of cooling power/over head) continued through the weekend.
  • HMS collimator issue was fixed.
  • HMS Q2 Alarms were disabled in the alarm handler. One was turned back on.
  • HMS Dipole Diff alarm was disabled (due to faulty readout and persistent alarm).
  • HMS door controller reset

Planned Accesses to the Hall

  • Will need to re-establish communication with the SHMS Q1 read-back current while changing to phase-2 kinematic.
  • Joe will fix something with the measuring line voltages.
  • Walter will test vacuum pumps tomorrow morning.
  • All SHMS magnets leads de ice. Tomorrow.

Planned Accelerator Configuration Changes

Special Requests

  • Latif's GEM testing completed this morning during beam studies.